Grade 7Magical Mathematics[Interesting Approach]

Write the equivalent rational numbers of -5/3 with numerator is -15

lakshminarayana appana
4 Years agoGrade 7

1 Answer

Yogendra Kumar
4 Years ago
Hi

the equivalent rational numbers of -5/3 with numerator is -15

\frac{-5}{3}=\frac{-5\times 3}{3\times3}=\frac{-15}{9}

Thanks and Regards